# Remote Work and Women’s Labor Supply: The New Gender Division at Home

### v1

@@ 6 Mechanisms
Table 3: Effect on Time Spent in Labor and Household Supply - Share of the Day
@@
$Exposed_{o(H)}\times Post_{t}$ 0.075*** 0.074*** 0.054*** 0.021 -0.000 0.000 -0.013 0.019*
@@
$Exposed_{o(H)}\times Post_{t}$ 0.024 0.023 0.047*** -0.024 0.005* -0.009** -0.001 -0.024*
@@
$Exposed_{o(H)}\times Post_{t}$ -0.001 -0.004 -0.006* -0.027** 0.022 0.025 -0.002 0.084***
@@
Mean pre-treatment 0.093 0.045 0.016 0.106 0.123 0.228 0.334 0.645
Obs 7,989 7,989 7,989 7,989 4,691 7,989 7,989 7,849
$R^{2}$ 0.86 0.83 0.85 0.84 0.92 0.84 0.84 0.85
@@
$Exposed_{o(H)}\times Post_{t}$ 0.003 0.002 -0.004* -0.004 0.028 0.001 -0.003 -0.049
@@
Mean pre-treatment 0.041 0.016 0.007 0.051 0.042 0.127 0.177 0.649
Obs 8,186 8,186 8,186 8,186 7,113 8,186 8,186 7,568
$R^{2}$ 0.85 0.81 0.84 0.86 0.87 0.83 0.83 0.86
@@
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. We show results for the time spent in labor, household and personal activities available in ATUS. Time-use variables are recorded by one member in the household. Then, as Section  2 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. While the former has data on the time spent doing activities by husbands, the latter will have for wives. There is a first panel that has dependent variables that capture time spent in labor production, as a share of the whole day: working, working at their main work, working at their main work from home and outside home, working at their secondary work and commuting to work. Those that do not work should have a 0 imputed in all working variables. There is a panel with dependent variables related to personal care (i.e. sleeping or taking a shower) and leisure (ie. going to religious events or sport activities). Lastly there is a household panel with dependent variables that capture the time spent doing the following activities: housework, cooking, shopping, primary childcaring (report doing childcaring as main activity), childcaring while working, total childcaring as secondary activity (childcaring while doing primary activities), total childcaring (the sum of primary and secondary childcaring) and the share of secondary childcaring on total childcaring. We restrict our sample to weekdays and non-holiday days. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the occupation level.
The substitution in household production between spouses varies with the age of the youngest child. Table B.31 shows, to a certain degree, that when the youngest child is no longer a toddler, women are more likely to work longer hours - particularly outside home - and reduce time spent in household tasks. Moreover, secondary childcare increases in relative importance.
On the other hand, husbands in these households exhibit a significant reallocation of where their main job takes place, without changing their total working time. Specifically, the likelihood of working from home increases by 8.7 percentage points, accompanied by a 10 percentage point decline in working outside the home. Consistent with this shift, husbands report more time spent on household chore –although not fully significant as the multiple hypothesis test suggests – and a significant 5.5 percentage point rise in total childcare time, largely driven by primary childcare, though the latter component is not statistically significant. [^34] Those magnitudes come from the sum of the main and interaction effects. Nonetheless, these adjustments do not fully offset the decline in time contributed by their wives. In addition, husbands spend significantly less time on childcare while working.
In contrast, households with toddlers exhibit a different pattern. Among women, there is no significant change in overall working time. However, the share of time spent working from home increases by 6.9 percentage points, while work outside the home declines by 4.7 percentage points, consistent with a reduction in commuting time. These women also allocate more time to household tasks such as shopping, but show limited changes in childcare, with the exception of a significant rise in time spent caring for children while engaged in other activities.
Among husbands in toddler households, total working time increases, driven by a shift towards working from home. They are also significantly less likely to hold a secondary job. In terms of household production, there is no significant change in time spent on most domestic tasks, except for cooking, which increases. As expected given the age of their youngest children, these husbands spend more time caring for them while working. However, they do not report significant changes in either primary or total childcare time.
To reinforce the validity of the findings in this section, we perform several robustness checks. We begin by evaluating the comparability between the two ATUS samples and the CPS, given that the ATUS samples are non-overlapping subsamples drawn from the CPS. Table B.32 presents the first and second moments of the distribution for key characteristics of husbands, wives, and households across the CPS sample and both ATUS subsamples. The two ATUS samples exhibit nearly identical means and standard deviations across all variables, with the exception that individuals in the men-respondent sample are more likely to be classified as treated. Additionally, there are differences in the second moment of wages between these subsamples. The CPS sample displays modest differences relative to the ATUS samples, particularly in the likelihood of treatment and in measures of household income and education. These patterns suggest that the ATUS subsamples tend to capture relatively wealthier and more educated households.
A further concern is that the estimated effects might be driven by specific subsamples whose responses to the WFH shock diverge from those identified in the main CPS analysis in Table 2. We address this in two ways. First, both Table 3 and Figure A.7 show that the increase in women’s employment following the WFH shock is statistically significant and persistent. Second, despite the smaller sample sizes in ATUS, Table B.33 shows that the estimated effects on wives’ labor market outcomes are directionally consistent with those in the main CPS sample.
These exercises yield two key implications. First, although the women- and men-respondent samples are drawn independently, their similarity implies that results based on one can be generalized to the other. In other words, if time-use diaries were available for both spouses within the same household, we would expect patterns similar to those observed across the separate respondent samples. Second, the alignment between the ATUS and CPS results suggests that the findings in the ATUS subsamples can, to some extent, be extrapolated to the broader CPS sample.
Finally, we conduct a standard placebo test by assigning the onset of the post-treatment period to March 2018. Table B.34 reports the results. We find no significant changes along all the set of variables, further supporting the credibility of our baseline results.
Childcare in SIPP
We are also able to test the mechanisms using the SIPP childcare topical modules, which are available for all years in our 2018-2024 SIPP panels, and ask reference parents several (extensive margin) question about any agents involved in childcare activities for their children.
We present some results validating both mechanisms in the top panel of Table 4. We estimate our baseline cross-section specification on the 2018-2024 SIPP panel datasets on a battery of childcare related variables. Our results show that in households shocked by the occupation-level WFH treatment, when the child’s mother is working, the spouse is 6.4 pp more likely to be involved in childcare, or roughly a 13% increase over the baseline. The woman herself is not any more likely to be involved in childcare by 1.1 pp while she is at work. This pair of results are most consistent with the substitution hypothesis.
We also find that households treated by the WFH shock are significantly less likely to recur to the child’s grandparent - or any other relatives - or a daycare for day-to-day childcare. Further, households are 7.3 pp less likely to have paid for childcare and, although not significant, overall expenditure in childcare decreased by some 16%. This is consistent with the finding that male parents are more available for childcare, once again, consistent with the substitution channel hypothesis.
To validate the identification strategy underpinning these past estimations, we run a placebo test analogous to Table B.24, where we estimate the same specification on the 2017-2019 SIPP subsample and using artificial timing for the start of treatment (March 2018). We present this placebo test on the SIPP childcare and child activities outcomes on Table B.35. The falsification tests are largely reassuring: we find no significant effects on parental involvement in childcare during the wife’s working hours, nor on household expenditures on childcare. However, as a caveat, we note that the estimated effects on childcare provided by grandparents and other relatives are negative.
@@
Table 4: Effect of Husbands’ WFH Occupation Shocks on Childcare Time and Child-Related Activities
@@ A. Entire sample
$Exposed_{o(H)}\times Post_{t}$ 0.004 0.064* -0.001 0.023 -0.009 -0.053* -0.073** -0.165

### v2

@@ C. Trim top 5% income earners
Table B.34: Effect of Husbands’ WFH Shock on Time Spent in Labor and Household Supply - Share of the Day
@@
$Exposed_{o(h(i))}\times Post_{t}$ 0.072*** 0.071*** 0.049*** 0.022 -0.000 0.001 -0.010 0.025**
(0.017) (0.017) (0.013) (0.016) (0.001) (0.002) (0.011) (0.011)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.080*** -0.083*** -0.039** -0.044** 0.002** -0.000 0.005 -0.030***
@@
$Exposed_{o(h(i))}\times Post_{t}$ 0.013 0.013 0.051*** -0.038** 0.004 -0.012*** 0.006 -0.023*
(0.016) (0.016) (0.011) (0.018) (0.003) (0.004) (0.011) (0.014)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.016 -0.012 0.049*** -0.061** -0.003 -0.000 0.023** 0.010
@@
$Exposed_{o(h(i))}\times Post_{t}$ -0.007 -0.006 -0.006** -0.025** 0.012 0.011 -0.014 0.060**
@@
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ 0.009 0.008 0.004 0.050*** -0.022** -0.005 0.045* -0.064**
(0.011) (0.006) (0.003) (0.013) (0.011) (0.022) (0.025) (0.032)
Mean pre-treatment 0.093 0.045 0.016 0.106 0.016 0.228 0.334 0.645
$\beta_{1}+\beta_{2}=0$ (p-val) 0.831 0.625 0.519 0.027 0.363 0.786 0.190 0.882
Obs 11,235 11,235 11,235 11,235 11,235 11,235 11,235 11,046
$R^{2}$ 0.84 0.85 0.82 0.82 0.83 0.82 0.82 0.83
@@
$Exposed_{o(h(i))}\times Post_{t}$ 0.001 0.002 -0.002 0.000 0.015** 0.011 0.012 -0.053
@@
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ 0.001 -0.006* 0.004 0.004 0.022 0.019 0.023 0.035
(0.011) (0.003) (0.003) (0.007) (0.013) (0.029) (0.029) (0.047)
Mean pre-treatment 0.041 0.016 0.007 0.051 0.008 0.127 0.177 0.649
$\beta_{1}+\beta_{2}=0$ (p-val) 0.811 0.125 0.318 0.531 0.004 0.223 0.142 0.713
Obs 11,440 11,440 11,440 11,440 11,440 11,440 11,440 10,585
$R^{2}$ 0.83 0.88 0.85 0.84 0.80 0.80 0.81 0.83
@@
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. We show results for the time spent in labor, household and personal activities available in ATUS. Time-use variables are recorded by one member in the household. Then, as Section 3 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. While the former has data on the time spent doing activities by husbands, the latter will have for wives. There is a first panel that has dependent variables that capture time spent in labor production, as a share of the whole day: working, working at their main work, working at their main work from home and outside home, working at their secondary work and commuting to work. Those that do not work should have a 0 imputed in all working variables. There is a panel with dependent variables related to personal care (i.e. sleeping or taking a shower) and leisure (ie. going to religious events or sport activities). Lastly there is a household panel with dependent variables that capture the time spent doing the following activities: housework, cooking, shopping, primary childcaring (report doing childcaring as main activity), childcaring while working, total childcaring as secondary activity (childcaring while doing primary activities), total childcaring (the sum of primary and secondary childcaring) and the share of secondary childcaring on total childcaring. We restrict our sample to weekdays and non-holiday days. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the occupation level.
Table B.35: Effects of Husbands’ Occupation and Women College Major WFH Shocks on Women WFH Uptake and Commuting
WFH Uptake Commuting
$Exposed_{o(h(i))}\times Post_{t}$ 0.019*** -0.373
(0.005) (0.265)
$Exposed_{e(i)}\times Post_{t}$ 0.044*** -2.287***
(0.008) (0.385)
$Exposed_{o(h(i))}\times Exposed_{e(i)}\times Post_{t}$ 0.038*** -0.704
(0.010) (0.464)
Mean pre-treatment 0.083 24.488
Obs 482,358 482,298
$R^{2}$ 0.11 0.04
  • Notes: This table presents estimates from Equation (7) in each column. Treatment is defined based on the change in the probability of WFH before and after the COVID-19 pandemic, using job postings data at the husband’s occupation ($o(h(i))$, Census 2010) levels and the actual WFH uptake data by women’s major college degree ($e(i)$). The sample is the 2017–2023 ACS sample of dual-parent households with children of child-rearing age. All regressions include fixed effects for state-by-time, husband’s occupation, and women’s major college degree. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
@@
Table B.36: Effect of Husbands’ WFH Occupation Shocks on Childcare Time and Child-Related Activities
@@ A. Entire sample
$Exposed_{o(h(i))}\times Post_{t}$ 0.004 0.064* -0.001 0.023 -0.009 -0.053* -0.073** -0.165